OpenStudy (anonymous):

Solve using an appropriate method. 5x -9y = 7, 7y -3x = 5

OpenStudy (anonymous):

rewrite so the x and y line up 5x-9y = 7 -3x+7y=5 now lets eliminate a variable multiple top by 3, bottom by 5 to cancel out 15x-27y=21 -15x+35y=25 add them together 8y=46 y = 46/8 you can do solve for x by plugging in y

OpenStudy (anonymous):

I guess Im not uderstanding how you get the 45x-27y=21 or the -15x=35y=25

OpenStudy (kira_yamato):

Did I do something wrong here?

OpenStudy (anonymous):

you want to change the 2 equations, so one of the variables cancel out, i picked x just out of preference. to do this you need to make the x in both equations cancel out

OpenStudy (anonymous):

can you show me the process of that because Im not totally getting it. I guess I need to see it broke down

OpenStudy (kira_yamato):

I'm using the matrix method but I got something different.... Integers in fact for both x and y... And they're huge...